Jean Poole is the personnel manager at Amgen. He recently reviewed the 300 applicants for several financial analyst positions Amgen has open. Sixty of the applicants hold MBA degrees. Each applicant was given a financial principles competency exam. One hundred thirty five passed the exam but were not MBA graduates, while 15 were MBA graduates who failed the exam.

a. What proportion of the applicants passed the exam?

b. What proportion of the applicants were MBA graduates who passed the exam? c. What proportion of the applicants were either MBA graduates or passed the exam? d. If the applicant is a MBA graduate, what is the probability of passing the exam? e. If the applicant is not a MBA graduate, what is the probability of passing the exam? f. Are the events independent? Mutually exclusive? Collectively exhaustive?
Answer #1
MBA not MBA total
passed 45 130 175
failed 15 110 125
total 60 240 300

a)

proportion of the applicants passed the exam =175/300=7/12 =0.5833

b)

proportion of the applicants were MBA graduates who passed the exam =45/300=0.15

c)

proportion of the applicants were either MBA graduates or passed the exam =(60+175-45)/300 =0.6333

d)

P(pass |MBA)=45/60 =0.75

e)

P(pass |not MBA)=130/240 =0.5417

f)

Collectively exhaustive